OpenStudy (anonymous):

PLEASE HELP The principal will randomly choose 6 students from a large school to represent the school in a newspaper photograph. The probability that a chosen student is an athlete is 30% (assume that this doesn't change). What is the probability that 4 athletes are chosen?

OpenStudy (anonymous):

The chance exactly 4 athletes are chosen implies the other 2 are none athletes. We use the binomial formula to calculate in how many different ways we can do this. We multiply those two and get the answer \[(0.3)^4 \times (0.7)^2 \times \left(\begin{matrix}6 \\ 4\end{matrix}\right) = 0,003969 \times 15 = 0,059535 \approx 6\% \]